How do you calculate the mass compound and molar concentration of NaCl, KCl, Na2HPO4, KH2PO4
1
Expert's answer
2020-07-02T07:05:50-0400

Molar mass is calculated by summing the mass of each compound

1. AW of Na: 23, P: 31, O: 16, H: 1)

Na2HPO4 = 46+1+31+4x16 = 142


2. KCl

AW of K: 39, Cl: 35.4

FW of KCl=39 + 35.4 =74.4

3. NaCl

(AW of Na: 23, Cl: 35.5)

FW of NaCl =23+35.45 = 58.5

4.KH2PO4

K:39, H:1,P:31,O:16

FW of KH2PO4 =39+2+31+16×4= 136

4. Molar concentration is mass concentration divided by molar mass